OpenStudy (anonymous):

derivative y=sqrt(x^2+6x*sqrt(x))

OpenStudy (dumbcow):

chain rule u = x^2 +6xsqrtx du = 2x +6sqrtx + 3x/sqrtx y' = du/2sqrt(u)

OpenStudy (anonymous):

u' = 2x + 6 sqrtx + 3 sqrtx = 2x + 9 sqrt x => y' = (2x + 9 sqrt x )/ 2 sqrt ( x^2 + 6x sqrtx)